OpenStudy (dboyette):

# of absences/final grade 0 91 1 88 2 85 3 78 4 66 5 51 Find the correlation coefficient for the data. Find the equation for the regression line for the data, and predict the final grade of a student who misses 3.5 days.

OpenStudy (zarkon):

r=-.950896325398 \[\hat{y}(x)=-7.8x+96\] \[\hat{y}(3.5)=-7.8*3.5+96=68.7\]